Exercise 16-08

The following are two independent situations.

1. Sandhill Cosmetics acquired 10% of the 275,500 shares of common stock of Nevins Fashion at a total cost of $15 per share on March 18, 2020. On June 30, Nevins declared and paid a $73,300 dividend. On December 31, Nevins reported net income of $128,800 for the year. At December 31, the market price of Nevins Fashion was $17 per share.
2. Blue, Inc., obtained significant influence over Rogan Corporation by buying 30% of Rogan’s 38,400 outstanding shares of common stock at a total cost of $8 per share on January 1, 2020. On June 15, Rogan declared and paid a cash dividend of $35,800. On December 31, Rogan reported a net income of $77,700 for the year.


Prepare all the necessary journal entries for 2020 for (a) Sandhill Cosmetics and (b) Blue, Inc. (Credit account titles are automatically indented when amount is entered. Do not indent manually. Record journal entries in the order presented in the problem. If no entry is required, select "No entry" for the account titles and enter 0 for the amounts.)

Date

Account Titles and Explanation

Debit

Credit

(a)

Jan. 1Mar. 18June 15June 30Dec. 31

Jan. 1Mar. 18June 15June 30Dec. 31

Jan. 1Mar. 18June 15June 30Dec. 31

(b)

Jan. 1Mar. 18June 15June 30Dec. 31

Jan. 1Mar. 18June 15June 30Dec. 31

Jan. 1Mar. 18June 15June 30Dec. 31
Answer #1

(a) journal entries in the books of Sandhill Cosmetics

Fair value method
Date General Journal Debit Credit
March 18 Investment - Nevins fashions $ 413250
Cash (275500*10%= 27550) (27550*15) $ 413250
(To record Purchase of Investment.)
June 30 Cash $ 7330
Dividend Income $ 7330
(To record Dividend Received.) (73300*10%)
December 31 No Journal entry Required for Net income.
December 31 Investment - Nevins fashions $ 55100
Unrealized Gain/Loss - other comprehensive income $ 55100
(To record Fair Value Adjustment.) ((17-15)*27550)

Sandhill Cosmetics acquired 10% shares of Nevins Fashion it means investment is less than 20% , when will investee co's investment is less than 20% significant intrest should not excersised . here we use fair value method the reason behind that sandhil have not significant interest in nevins fashion so fair value method is used.

b) journal entries in the books of Blue, Inc.

Equity method
Date General Journal Debit Credit
January 1 Equity Investment In Rogan $ 92160
Cash (38400*30%= 11520) (11520*8) $ 92160
(To record Purchase of Investment.)
June 15 Cash $ 10740
Equity Investment In Rogan $ 10740
(To record Dividend Received.) (35800*30%)
Dec 31 Equity Investment In Rogan $ 23310
Income from Equity Investment (77700*30%) $ 23310
(To record Income from Equity Investment. )
Dec 31 No Journal entry Required for Fair Value Adjustment

blue inc acquired 30% shares of Rogan’s it means investment is more than 20% , when will investee co's investment is more than 20% significant interest should excersised . here we use Equity method the reason behind that blue inc  have significant interest in rogans  so equity method is used.